OpenStudy (anonymous):

Which of the following is a polynomial with roots 3, -3i, and 3i ? x3 + 3x2 + 9x + 27 x3 - 3x2 + 9x - 27 x3 + 9x2 + 9x + 81 x3 - 9x2 + 9x - 81 The little "i" thingies confuse me. :c

OpenStudy (anonymous):

i=sqrt(-1) i^2 =-1 i^3=(i)(i^2)=-i polynomial with 3, -3i, 3i is (x-3)(x+3i)(x-3i) =(x^2+3ix-3x-9i)(x-3i) =x^3-3ix^2+3ix^2+9x-3x^2+9ix-9ix-27 =x^3-3x^2+9x-27

OpenStudy (anonymous):

Thank you sooo much :3 You're fantastic.

OpenStudy (anonymous):

You are welcome